OpenStudy (faithspring):

solve by factoring 2. 4x2 + 23x – 6 = 0

OpenStudy (anonymous):

4x^23x-6 factored is: (4x-1)(x+6) 4x-1=6 4x=7 x= 7/4 x+6=0 x= -6

OpenStudy (faithspring):

how did you get the -6 and the 7? @ijakez

OpenStudy (faithspring):

ohhhhh

OpenStudy (faithspring):

nevermind

OpenStudy (math&ing001):

This equation has 2 solutions, one is -6 but the other isn't 7/4. 4x-1=0 => x= 1/4

OpenStudy (faithspring):

so the answer would be -6, and 1/4? @@math&ing001

OpenStudy (anonymous):

oh oops I mistook my own 1 for a 7 Yeah it'd be -6 and 1/4 sorry

OpenStudy (anonymous):

"4x-1=6" is actually supposed to be: 4x-1=0 sorry I just woke up add 1 to both sides 4x=1 divide both sides by x x=1/4

OpenStudy (math&ing001):

Yes correct.

OpenStudy (faithspring):

ohhh okay. could you guys help with a couple more?

OpenStudy (math&ing001):

I have to get off for now, but I'll be back in about an hour.

OpenStudy (faithspring):

hmm okay thank you

OpenStudy (anonymous):

I can help you after I grab some tea to wake up

OpenStudy (faithspring):

okay thank you, I'll type up a few more questions while you get tea @ijakez

OpenStudy (faithspring):

(still factoring) 1. x^2 +7+10=0 3.x^2-16=0 4.2x^3 +10x^2-12x+0 5. 24x^2= 2 - 13x ( you dont have to answer all of them, you can choose the ones you want to answer )

OpenStudy (anonymous):

I just typed up the answers for 3 of them and then realized how stupid that is. Do you know how to do these yourself?

OpenStudy (faithspring):

Sorta, after you guys explained how to do it, it kinda clicked but at the same time I still don't quite get it.

OpenStudy (faithspring):

@ijakez also for number 2, wouldn't it be -1/4

OpenStudy (anonymous):

Watch this video. The second problem is not -1/4. https://www.youtube.com/watch?v=1kfq0aR3ASs

OpenStudy (faithspring):

so I solved 1 and I watched the video

OpenStudy (faithspring):

but could you maybe help with 3 and 5? @ijakez